OpenStudy (saifoo.khan):

In a box of tickets,there are 50 blue tickets and 15 red tickets.after a certain number of red tickets are added to the box,the probability of randomly choosing a red ticket is 2/3.how many red tickets were added to the box?

OpenStudy (saifoo.khan):

@satellite73

OpenStudy (anonymous):

i bet you can do this

OpenStudy (saifoo.khan):

I forgot.

OpenStudy (saifoo.khan):

I know we will make something equal to 2/3 and suppose the cards as x.

OpenStudy (anonymous):

add \(x\) red ones so you have a total of \(x+15\) red and \(65+x\) in the box ratio of red to total is \[\frac{15+x}{65+x}=\frac{2}{3}\]

OpenStudy (saifoo.khan):

@satellite73 is already helping, thanks.

Parth (parthkohli):

Okay. Sorry to interfere :)

OpenStudy (saifoo.khan):

I got my memory back! :D

OpenStudy (saifoo.khan):

number/TOTAL !!

OpenStudy (anonymous):

whew! you had me worried for a second

OpenStudy (saifoo.khan):

;D